OpenStudy (hippoharry):

Solve the equation below by completing the square and then solving for x. x^2 + 14x + 24 = 0 A. x = 8 or x = 3 B. x = -12 or x = -2 C. x = 2 or x = -2 D. x = -2 or x = -5

OpenStudy (hippoharry):

@Erak do i divide by 14?

OpenStudy (erak):

Take 14/2, then square that.

OpenStudy (hippoharry):

i got seven. do i then square that?

OpenStudy (erak):

Yeah then you do something where you add and subtract that number, like so: x^2 + 14x + 49 + 24 - 49 = 0 Then move the original 24 and the -49 over: x^2 + 14x + 49 = -24 + 49 Recognise the left side turns into a perfect square and simplify the right (x+7)^2 = 25 Try the rest

OpenStudy (hippoharry):

x=-12 and x=-2

OpenStudy (erak):

yeah, because you get +5 and -5 when you square root the 25, and 7 subtracted from those numbers are -12 and -2